OpenStudy (anonymous):

Find the perimeter of rectangle with the given base and height. 21in., 7in Find the perimeter of rectangle with the given base and height. 21in., 7in @Mathematics

OpenStudy (hoblos):

56 in

OpenStudy (anonymous):

\[2l+2h\]\[=(2\times 21) + (2\times 7)\]\[=42+14\]\[=56\]

OpenStudy (anonymous):

Thank you!